# DEFT: Differentiable Branched Discrete Elastic Rods for Modeling Furcated DLOs in Real-Time

### v1

@@ Abstract
Autonomous wire harness assembly requires robots to manipulate complex branched cables with high precision and reliability. A key challenge in automating this process is predicting how these flexible and branched structures behave under manipulation. Refer to caption Figure 1: This paper introduces DEFT, a novel framework that combines a differentiable physics-based model with a learning-based approach to accurately model and predict the dynamic behavior of Branched Deformable Linear Objects (BDLOs) in real time. As this paper illustrates, this model can be used in concert with a motion planning algorithm to autonomously manipulate BDLOs. The figures above illustrate how DEFT can be used to autonomously complete a wire insertion task. Top: The system first plans a shape-matching motion, transitioning the BDLO from its initial configuration to the target shape (contoured with yellow), which serves as an intermediate waypoint. Bottom: Starting from the intermediate configuration, the system performs thread insertion, guiding the BDLO into the target hole while also matching the target shape. Notably, DEFT predicts the state of the wire recursively without relying on ground truth or perception data at any point in the process. Without accurate predictions, it is difficult for robots to reliably plan or execute assembly operations. While existing research has made progress in modeling single-threaded Deformable Linear Objects (DLOs), extending these approaches to Branched Deformable Linear Objects (BDLOs) presents fundamental challenges. The junction points in BDLOs create complex force interactions and strain propagation patterns that cannot be adequately captured by simply connecting multiple single-DLO models. To address these challenges, this paper presents Differentiable discrete branched Elastic rods for modeling Furcated DLOs in real-Time (DEFT), a novel framework that combines a differentiable physics-based model with a learning framework to: 1) accurately model BDLO dynamics, including dynamic propagation at junction points and grasping in the middle of a BDLO, 2) achieve efficient computation for real-time inference, and 3) enable planning to demonstrate dexterous BDLO manipulation. A comprehensive series of real-world experiments demonstrates DEFT’s efficacy in terms of accuracy, computational speed, and generalizability compared to state-of-the-art alternatives.
I Introduction
@@
The automation of wire harness assembly represents a critical challenge in manufacturing, particularly in the automotive and aerospace industries where complex, branched cable systems are essential components. These assemblies require robots to manipulate Branched Deformable Linear Objects (BDLOs) with precision over extended periods ($\geq$5s) [1, 2, 3], a task that remains largely unautomated due to the objects’ complex nonlinear dynamics. While accurate prediction of BDLO behavior is possible through computationally intensive models, real-time manipulation demands rapid computation. Further complicating matters, practical assembly environments often involve occlusions that prevent perception systems from reliably estimating the full BDLO configuration.
Existing approaches to modeling BDLOs  [4, 5, 6, 7] have focused primarily on computer graphics applications, prioritizing visual plausibility over physical accuracy. These models typically require extensive manual parameter tuning and make impractical assumptions about perfect state observation—limitations that render them unsuitable for robotic manipulation tasks. Recent advances in modeling single Deformable Linear Objects (DLOs) [8, 9, 10, 11, 12] have shown promise, but the extension to branched structures introduces fundamental new challenges. When a robot manipulates one branch of a wire harness, the forces and deformations propagate through junction points to affect all connected branches. This coupling between branches, combined with the need for flexible grasp locations and real-time computation, makes BDLO manipulation significantly more complex than handling single DLOs.
To address these challenges, this paper introduces Differentiable Branched Discrete Elastic Rods for Modeling Furcated DLOs in Real-Time (DEFT). This framework introduces a novel junction coupling approach that accurately models dynamic propagation between parent and child branches while maintaining real-time performance. As this paper illustrates through real-world evaluations, DEFT enables three essential wire harness assembly tasks: 3D shape matching, multi-thread grasp, and precise thread grasping and insertion. The contributions of this paper are four-fold:
  1. 1.
    A novel differentiable BDLO model that accurately captures how dynamics propagate through wire harness junction points thereby enabling model parameters to be efficiently identified via experimental observation;
  2. 2.
    A computational representation that enables efficient inference and model parameter learning through parallel programming in concert with analytical gradients;
  3. 3.
    A specialized Graph Neural Network (GNN) architecture for BDLO residual learning, which can compensate for numerical integration errors that arise over long time horizon prediction;
  4. 4.
    A comprehensive experimental validation demonstrating DEFT’s superior performance in accuracy, speed, and sampling efficiency compared to existing approaches, including integration into a motion planning framework for real-world manipulation tasks.
The framework is implemented in PyTorch for seamless integration with deep learning workflows, and represents, to the best of our knowledge, the first quantitative investigation of BDLO modeling and manipulation in real-world settings. The results demonstrate that DEFT enables reliable autonomous manipulation of complex wire harnesses, addressing a significant challenge in manufacturing automation. The code and dataset associated with this paper can be found at https://github.com/roahmlab/DEFT.
II Related Work
Effective BDLO modeling requires three critical capabilities: real-time prediction, accurate 3D simulation, and efficient model parameter learning. Real-time prediction capability is essential for robotic manipulation tasks, where quick decision-making can mean the difference between successful and failed object handling. Three-dimensional accuracy ensures the model can capture the complex spatial deformations that BDLOs undergo during manipulation, including twisting, bending, and branch interactions. Efficient parameter learning enables models to automatically adapt to different materials and configurations through data-driven optimization, reducing both computational overhead and the need for extensive training data. This section reviews relevant approaches in modeling BDLOs and single DLOs while considering each of these criteria.
II-A DLOs Modeling
Approaches to modeling DLOs have evolved from purely physics-based methods toward more sophisticated learning-based frameworks that balance between the three aforemention key capabilities.
Physics-based approaches implement different mathematical frameworks to model physical behavior. Mass-spring systems represent objects as networks of point masses connected by springs, computing forces and accelerations to simulate movement [13, 14, 15]. Position-Based Dynamics (PBD) directly manipulates vertex positions to satisfy physical constraints, avoiding expensive force calculations [10, 16, 9]. Finite Element Methods (FEM) divide objects into discrete elements and solve partial differential equations to compute deformations [17, 18, 19]. Discrete Elastic Rods (DER) models represent objects as centerline curves with material frames, enabling explicit modeling of bending and twisting [16, 20, 21, 22]. Unfortunately each of these methods has limitations: mass-spring systems describe the rigidity of wires by introducing large spring stiffness which can make prediction numerically unstable; PBD models struggle with parameter learning because they assume that the system is quasi-static; FEM methods are computationally intensive; and DER models lack differentiable expressions for their model parameters, making efficient parameter optimization challenging.
Learning-based modeling approaches use data-driven techniques to predict deformation behavior. Methods such as Graph Neural Networks (GNN) process vertex-level spatial information [12], while Bi-LSTM models capture temporal dependencies in deformation sequences [11]. These approaches excel at real-time prediction through efficient computation and support straightforward parameter learning through standard optimization techniques. However, they struggle with 3D accuracy, particularly in capturing the complex dynamic deformations characteristic of real-world manipulation.
Recent work has explored hybrid approaches that combine learning with physics-based modeling. This is exemplified by DEFORM [8] which constructs a novel differentiable DER model and combines it with a neural network-based residual learning approach to account for numerical integration errors during prediction. This hybrid approach achieves both real-time performance and 3D accuracy while enabling efficient parameter learning through end-to-end differentiability.
Despite DEFORM’s advances in DLO modeling, it faces three significant limitations when extended to BDLO applications. First, DEFORM’s modeling framework does not address the dynamics of branched structures, particularly the propagation of stretching, twisting, and bending forces at junction points. Second, DEFORM’s manipulation framework assumes fixed clamping points at the object ends, following conventional DLO literature. This assumption restricts its applicability to more sophisticated manipulation tasks, such as thread insertion, which require grasping at multiple points along the object’s length. Third, DEFORM’s computational approach presents efficiency challenges for BDLO simulation. Its sequential processing of individual branches would result in linear computational time increases as branches are added. Furthermore, its bi-level optimization framework, while enabling learning from real-world data, relies on numerical gradients that can lead to slow convergence and optimization instability.
II-B BDLOs Modeling
Existing BDLO modeling approaches adapt various physical simulation frameworks to handle branched structures. FEM-based methods extend traditional finite element analysis by treating branches as connected elastic elements [23]. While these methods excel in 3D accuracy and enable efficient parameter learning through gradient-based optimization, their computational requirements preclude real-time prediction. DER-based approaches introduce rigid-body constraints at junctions to ensure proper alignment of attached segments  [4], achieving real-time performance and reasonable 3D accuracy but struggling with parameter learning due to their fixed parameter structure. Quasi-static methods like ASMC enforce geometric constraints while assuming static equilibrium conditions [5], offering real-time performance, but sacrificing 3D accuracy by overlooking dynamic effects that can make parameter learning challenging.
To address these challenges, DEFT introduces a novel differentiable framework specifically designed for BDLOs. Through a specialized GNN architecture for branched structures, DEFT achieves all three critical capabilities: real-time performance, 3D accuracy, and efficient parameter learning. This enables DEFT to adapt to various materials and configurations while maintaining the speed and accuracy necessary for practical robotic manipulation tasks.
III Preliminaries
This section introduces the key components involved in modeling DLOs. It then summarizes the primary contributions of DEFORM and describes the challenges of adapting it to BDLO modeling.
III-A Deformable Linear Object (DLO) Model

### v5

@@ Abstract
Autonomous wire harness assembly requires robots to manipulate complex branched cables with high precision and reliability. A key challenge in automating this process is accurately predicting how flexible branched structures behave during manipulation. Without accurate predictions, robots cannot reliably execute assembly operations. While existing research has made progress in modeling single-threaded Deformable Linear Objects (DLOs), extending these approaches to Branched Deformable Linear Objects (BDLOs) presents fundamental challenges. The junction points in BDLOs create complex force interactions and strain propagation patterns that cannot be adequately captured by simply connecting multiple single-DLO models. To address these challenges, this paper introduces Differentiable discrete branched Elastic rods for modeling Furcated DLOs in real-Time (DEFT), a novel framework that combines a differentiable physics-based model with a learning framework to: 1) accurately model BDLO dynamics, including dynamic propagation at junction points and grasping in the middle of a BDLO, 2) achieve efficient computation for real-time inference, and 3) enable planning to demonstrate dexterous BDLO manipulation. A comprehensive series of real-world experiments demonstrate that DEFT outperforms state-of-the-art methods in accuracy, computational speed, and planning success rate.
Keywords: Branched Deformable Linear Objects Modeling, Physics-Informed Learning, Differentiable Simulation
Figure 1: As this paper illustrates, DEFT can be used in concert with a motion planning algorithm to autonomously manipulate BDLOs. The figures above illustrate how DEFT can be used to autonomously complete a wire insertion task. Left: The system first plans a shape-matching motion, transitioning the BDLO from its initial configuration to the target shape (contoured with yellow), which serves as an intermediate waypoint. Right: Starting from the intermediate configuration, the system performs thread insertion, guiding the BDLO into the target hole while also matching the target shape. Notably, DEFT predicts the state of the wire recursively without relying on ground truth or perception data at any point in the process.
1 Introduction
@@
The automation of wire harness assembly represents a critical challenge in manufacturing, particularly in the automotive and aerospace industries where complex, branched cable systems are essential components. These assemblies require robots to manipulate Branched Deformable Linear Objects (BDLOs) with precision over extended periods ($\geq$5s) [1, 2, 3], a task that remains largely unautomated due to the objects’ complex nonlinear dynamics. Further complicating matters, practical assembly environments often involve occlusions, limiting perception systems’ ability to estimate full BDLO configurations reliably [4]. Thus, robust assembly requires accurate BDLO models for reliable long-term open-loop predictions. While accurate prediction of BDLO behavior is possible through computationally intensive models, real-time manipulation demands rapid computation.
Deep neural networks (DNNs) for tree-structured data [5, 6, 7, 8] can achieve real-time prediction performance for BDLOs but often require large datasets and lack robustness to variations in BDLOs. A promising alternative combines differentiable physics models as high-fidelity priors with DNNs to improve sample efficiency and generalizability [9, 10, 11, 12, 13, 14, 15]. However, this approach has not yet been applied to BDLO dynamics due to two significant challenges: First, the field lacks an accurate, efficient, differentiable physics model for BDLOs. Current models compromise speed for accuracy and do not preserve gradients needed for parameter optimization. Second, to the best of our knowledge, there is no existing research demonstrating that DNNs can effectively learn residual dynamics specifically for BDLOs.
To address these challenges, this paper introduces Differentiable Branched Discrete Elastic Rods for Modeling Furcated DLOs in Real-Time (DEFT), a novel differentiable physics-based framework that accurately predicts dynamic BDLO behavior over long time horizons with real-time inference capabilities. The contributions of this paper are four-fold: First, a novel differentiable BDLO model that accurately captures dynamics propagation at junctions, enabling efficient parameter identification. Second, a computational representation that enables efficient inference and model parameter learning through parallel programming combined with analytical gradients. Third, a specialized Graph Neural Network (GNN) architecture for residual learning to correct numerical integration errors over long horizons. Fourth, a comprehensive experimental validation demonstrating that DEFT outperforms prior methods in accuracy, speed, and sample efficiency while enabling practical applications such as 3‑D shape matching, multi‑branch grasping, and thread insertion (Figure 1). The framework is implemented in PyTorch for seamless integration with deep learning workflows, and represents, to the best of our knowledge, the first quantitative study of BDLO modeling and manipulation in real-world settings. A visualization of DEFT’s contributions can be found in Appendix B.1.
2 Related Work
Physics-Based Modeling. Several methods adapt tree or cloth modeling to branched structures. FEM-based approaches [16, 17, 18, 19, 20, 21, 22] offer high 3D accuracy but require extensive parameter tuning and are computationally intensive. Discrete Elastic Rods (DER) [23, 24] model branches individually using elastic rod theory with rigid-body constraints at junction points. These methods are reasonably accurate, but rely on fixed parameters that limit learning and require precise orientation tracking, which is impractical for robotic tasks. Quasi-static methods like ASMC [25] prioritize speed by enforcing geometric constraints, which can make parameter identification difficult and compromise dynamic fidelity.
Learning-Based Modeling. Unlike physics-based approaches, neural networks learn from real-world data. While Tree-LSTM [5, 26] and Graph Neural Networks (GNNs) [6, 8, 7] can process vertex-level spatial information for BDLOs, our experiments show they struggle to accurately capture complex 3D dynamic deformations during actual manipulation tasks.
Physics-Learning Hybrid Modeling. Recent research combines physics-based modeling with learning [10, 11, 12, 13, 14, 15]. For instance, DEFORM [14] focuses on modeling single-threaded DLOs using a differentiable DER (DDER) model and residual learning to correct numerical integration errors. However, DEFORM lacks fundamental capabilities essential for accurate and efficient BDLO modeling. Our experiments show that extending DEFORM to BDLOs reduces computational efficiency and fails to capture branching point dynamics, leading to inaccurate predictions. DEFT addresses these limitations by adapting DDER for individual branches and introduces several novel contributions (as described in the Introduction) for accurate, efficient BDLO simulation. Comprehensive ablation studies and comparisons validate DEFT’s advantages over DEFORM extensions and state-of-the-art methods.
3 Preliminaries